# Quantum-secure multiparty deep learning

### v1

@@ 7.3 Detailed security analyses
First, we express the setting of our problem in the continuous-variable quantum key distribution (CVQKD) framework. In a standard CVQKD setting, two legitimate parties Alice and Bob wish to communicate a message under possible attack by a third-party eavesdropper Eve external to both of them. In the CVQKD protocol, normal operation results in Bob receiving the message from Alice in its entirety, and the mutual information between Alice and Eve (along with Eve’s Holevo information) must be bounded below an arbitrary threshold.
@@
The eavesdropper may operate under different models based on different capabilities; here, we restrict our analysis to individual attacks. Under individual attacks, Eve performs independent and identically distributed (i.i.d.) attacks on all incoming modes; that is, they prepare separable ancilla states that interact individually with one of the signal modes in the quantum channel. The ancilla states are then stored in a quantum memory until measurement and are measured independently of each other. Extensions to collective and coherent attacks are left to future work.
@@ 7.3.1 Weights Leakage
Here, $\sigma_{z}$ is the Pauli matrix and $V$ represents the variance of the quadrature operators. The upper and lower diagonal $2\times 2$ blocks of $\Sigma_{\text{TMSVS}}$ are the covariance matrices of the quadrature operators of Alice and Bob respectively, while the off-diagonal block matrices are the covariances between the quadratures of Alice and those of Bob. The variance $V$ is related to the mean photon number per pulse by the relation $\mu=\frac{1}{2}(V-1)$. This variance is the sum of Alice’s actual modulation variance of her two quadrature components and the vacuum shot-noise variance.
@@
The Holevo theorem provides an upper bound on the accessible information that can be extracted from a quantum system. Tracing out Eve’s subspace from the joint state |$\ket{\psi}\bra{\psi}$ | yields Alice’s and Bob’s joint substate, i.e., $\mathrm{Tr}_{E}(\ket{\psi}\bra{\psi})=\rho_{AB}$. Mathematically, the Holevo bound $\chi$ for the information held by Eve’s substate $\rho_{E}=\mathrm{Tr}_{AB}(\ket{\psi}\bra{\psi})$ about the quantum state $\rho_{AB}$ shared between Alice and Bob is given by [46]:
@@
Then, we have $S_{AB}=g(\nu_{1})+g(\nu_{2})$. Note that in Section 4 we presented these quantities using the occupation $\mu=\frac{V-1}{2}$ and $T=1$. We analyze the effect of optical loss where the transmission $T<1$ in Section 5.
@@ 7.3.3 Data leakage
In this scenario, the client operates legitimately and the server acts maliciously. Under our eavesdropper model, we assume that the server can perform any arbitrary individual attack upon receiving the verification state $\rho_{v}$. In particular, the server could be equipped to make quantum measurements, observe quantum correlations, and transmit non-Gaussian states.
@@ 8 Acknowledgments
The authors kindly thank Saumil Bandyopadhyay, Leshem Choshen and Rom Dudkiewicz for fruitful discussions and suggestions. K. Sulimany acknowledges the support of the Israeli Council for Higher Education and the Zuckerman STEM Leadership Program.

### v2

@@ 7.3 Detailed security analyses
First, we express the setting of our problem in the continuous-variable quantum key distribution (CVQKD) framework. In a standard CVQKD setting, two legitimate parties Alice and Bob wish to communicate a message under possible attack by a third-party eavesdropper Eve external to both of them. In the CVQKD protocol, normal operation results in Bob receiving the message from Alice, and the mutual information between Alice and Eve (along with Eve’s Holevo information) must be bounded below an arbitrary threshold.
@@
The eavesdropper may operate under different models based on different capabilities; here, we restrict our analysis to individual attacks. Under individual attacks, Eve performs independent and identically distributed (i.i.d.) attacks on all incoming modes; that is, they prepare separable ancilla states that interact individually with one of the signal modes in the quantum channel. The ancilla states are then stored in a quantum memory until measurement and are measured independently of each other. Extensions to collective and coherent attacks are left for future work.
@@ 7.3.1 Weights Leakage
Here, $\sigma_{z}$ is the Pauli matrix, and $V$ represents the variance of the quadrature operators. The upper and lower diagonal $2\times 2$ blocks of $\Sigma_{\text{TMSVS}}$ are the covariance matrices of the quadrature operators of Alice and Bob respectively, while the off-diagonal block matrices are the covariances between the quadratures of Alice and those of Bob. The variance $V$ is related to the mean photon number per pulse by the relation $\mu=\frac{1}{2}(V-1)$. This variance is the sum of Alice’s actual modulation variance of her two quadrature components and the vacuum shot-noise variance.
@@
The Holevo theorem provides an upper bound on the accessible information that can be extracted from a quantum system. Tracing out Eve’s subspace from the joint state |$\ket{\psi}\bra{\psi}$ | yields Alice’s and Bob’s joint substate, that is, $\mathrm{Tr}_{E}(\ket{\psi}\bra{\psi})=\rho_{AB}$. Mathematically, the Holevo bound $\chi$ for the information held by Eve’s substate $\rho_{E}=\mathrm{Tr}_{AB}(\ket{\psi}\bra{\psi})$ about the quantum state $\rho_{AB}$ shared between Alice and Bob is given by [50]:
@@
Then, we have $S_{AB}=g(\nu_{1})+g(\nu_{2})$. Note that in Section 4 we presented these quantities using occupations $\mu=\frac{V-1}{2}$ and $T=1$. We analyze the effect of optical loss where the transmission $T<1$ in Section 5.
@@ 7.3.3 Data leakage
In this scenario, the client operates legitimately, and the server acts maliciously. Under our eavesdropper model, we assume that the server can perform any arbitrary individual attack upon receiving the verification state $\rho_{v}$. In particular, the server could be equipped to make quantum measurements, observe quantum correlations, and transmit non-Gaussian states.
@@ 7.3.4 Data Leakage under quantum operations
8 Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
9 Code availability
The code used in this study is available from the corresponding author upon reasonable request.
10 Author information
10.1 Contributions
K.S. proposed the protocol, performed the security analyses and numerical simulations, and wrote the manuscript under the guidance of D.E. S.K.V. designed the PyTorch package for optical deep neural networks 7.2. R.H. assisted with the optical realization of the protocol and its quantum description 7.1. P.I. assisted in the quantum calculation of the Cramér-Rao bound 7.3.4. All authors contributed to the writing.
11 Ethics declarations
11.1 Competing interests
K.S and D.E. have filed a US Patent application 63677972 on quantum-secure multiparty deep learning. The other authors declare no competing interests.
@@ 12 Acknowledgments
The authors kindly thank Saumil Bandyopadhyay, Leshem Choshen and Rom Dudkiewicz for fruitful discussions and suggestions. K.S acknowledges the support of the Israeli Council for Higher Education and the Zuckerman STEM Leadership Program.